# Object-Centric Diffusion for Efficient Video Editing

### v1

@@ Abstract
Diffusion-based video editing have reached impressive quality and can transform either the global style, local structure, and attributes of given video inputs, following textual edit prompts. However, such solutions typically incur heavy memory and computational costs to generate temporally-coherent frames, either in the form of diffusion inversion and/or cross-frame attention. In this paper, we conduct an analysis of such inefficiencies, and suggest simple yet effective modifications that allow significant speed-ups whilst maintaining quality. Moreover, we introduce Object-Centric Diffusion, coined as OCD, to further reduce latency by allocating computations more towards foreground edited regions that are arguably more important for perceptual quality. We achieve this by two novel proposals: i) Object-Centric Sampling, decoupling the diffusion steps spent on salient regions or background, allocating most of the model capacity to the former, and ii) Object-Centric 3D Token Merging, which reduces cost of cross-frame attention by fusing redundant tokens in unimportant background regions. Both techniques are readily applicable to a given video editing model without retraining, and can drastically reduce its memory and computational cost. We evaluate our proposals on inversion-based and control-signal-based editing pipelines, and show a latency reduction up to 10 for a comparable synthesis quality.
1 Introduction
@@
Diffusion models [43, 12] stand as the fundamental pillar of contemporary generative AI approaches [38, 14, 19]. Their success primarily stems from their unparalleled diversity and synthesis quality, surpassing the capabilities of the earlier versions of generative models [20, 9]. On top of that, recent methods such as Latent Diffusion Models (LDMs) [38] exhibit scalability to high-resolution inputs and adaptability to different domains without requiring retraining or finetuning, especially when trained on large-scale datasets [41]. These traits catalyzed impactful adaptations of pretrained LDMs across a spectrum of applications, including text-guided image generation [38], editing [1], inpainting [38], as well as video generation [2] and editing [46, 33, 49].
@@
Nevertheless, diffusion models come with various trade-offs. One immediate drawback is their inefficiency due to the sampling process, which involves iterating a denoising neural network across numerous diffusion steps. Despite the availability of techniques like step distillation [29, 40] and accelerated samplers [26, 27, 43] that expedite image synthesis, efficient sampling solutions for video generation are still lacking. Besides, special attention must be paid to maintaining temporal coherency among frames when creating or modifying a video. Otherwise, it may result in flickering artifacts or a complete lack of correlation between frames. To tackle this challenge, approaches like diffusion inversion [43, 33] and cross-frame self-attention [33, 49] have been introduced, but all at the cost of significantly increasing the computational load.
@@
This paper centers on video editing models and presents novel solutions to enhance their efficiency. We first examine the current video editing frameworks, identifying the key elements that impede their effectiveness. These encompass memory overheads, such as those associated with attention-based guidance from diffusion inversion, as well as computational bottlenecks, including excessive cross-frame attention and an unnecessarily high number of sampling steps. We show that significant improvements can be achieved by adopting off-the-shelf optimizations namely efficient samplers and leveraging token reduction techniques in attention layers such as token merging (ToMe) [4, 3].
@@
Additionally, we argue that during video editing, users may be particularly sensitive to the quality of edited foreground objects compared to any potential degradations in the background regions. Consequently, we propose two new and efficient techniques that harness this object-centric aspect of editing applications. Our first solution, Object-Centric Sampling, involves separating the diffusion process between edited objects and background regions. This strategy enables the model to focus the majority of its diffusion steps on the foreground areas. Consequently, it can generate unedited regions with considerably fewer steps, all while ensuring faithful reconstructions in all areas. With our second solution, Object-Centric 3D ToMe, we incentivise each token reduction layer to merge more tokens within less crucial background areas and exploit temporal redundancies that are abundant in videos. The combination of such techniques, coined Object-Centric Diffusion (OCD), can seamlessly be applied to any video editing method without retraining or finetuning.
@@
As we demonstrate in extensive experiments, the combined implementation of these object-centric solutions enhances the quality of video editing in both salient foreground object and background, while considerably reducing the overall generation cost. We apply our proposed solutions to inversion-based and ControlNet-based video editing models, speeding them up by a factor of 10 and 6 respectively while also decreasing memory consumption up to 17 for a comparable quality (see Fig. 1).
@@
We introduce Object-Centric 3D ToMe, which reduces number of cross-frame attention tokens by fusing redundant ones predominantly in background regions.
@@ 2 Related work
There has been a surge in text-to-video generation methods using diffusion models. The early works inflate the 2D image generation architectures to 3D by replacing most operations, i.e. convolutions and transformers, with their spatio-temporal 3D counterparts [15, 13, 16, 42, 50]. Despite their high temporal consistency, these substantial model modifications requires extensive model training on large collection of captioned videos.
To avoid extensive model trainings, recent works adopts off-the-shelf image generation models for video editing in one-shot and zero-shot settings. One-shot methods rely on test-time model adaption on the editing sample which is still too slow for real-time applications [46, 6]. Zero-shot methods integrate training-free techniques into the image generation models to ensure temporal consistency across frames [33, 49, 17, 8, 24] most commonly by: i) providing strong structural conditioning from the input video [7, 49, 5] inspired by ControlNet [48]. ii) injecting the activations and attention maps extracted by diffusion inversion [43, 44] of the input video into the generation [33, 8]. iii) replacing the self-attention operations in the denoising UNet with cross-attention where the key and value tokens are being extracted from neighboring frames [46, 33, 49]. Introducing cross-frame attention greatly increases the temporal consistency especially when using more neighboring frames [49]. Despite their effectiveness, all these temporal solutions come with additional computational cost which is under explored in the literature.
@@
Due to their sequential sampling, diffusion models are computationally very expensive. Several studies analyze the denoising U-Net architecture to enhance its runtime efficiency, achieved either through model distillation [18] or by eliminating redundant modules [23]. Other ways to obtain efficiency gains include enhanced noise schedulers [26, 27] and step distillation techniques [40, 23, 29]. The most relevant to our work is the Token Merging technique presented in [4] that demonstrated advantages for image generation [3]. Although very proficient, we observe it does not readily work when deployed in video editing settings. For this reason, we optimize this technique for the task at hand, such that it exploits redundancy of tokens across frames and directing its token fusing towards background regions. Similar to our object centric sampling, [22] aims for an efficient editing by skipping the computations on unedited regions for interactive image editing: the user interactively edits different regions of an image. They cache the representation during generations to use it for unedited regions in future edits. Instead of caching the features, our proposed object centric sampling is not limited to interactive editing applications.
@@ 3 Efficiency Bottlenecks
To investigate the main latency bottlenecks in video editing pipelines, we run a benchmark analysis on a representative pipeline, derived from FateZero [33]. More specifically, in this case-study we are primarily interested in the role of components such as diffusion inversion and cross-frame attention, as these techniques are crucial for the task and ubiquitous in literature [46, 49, 33, 8]. To study the impact of such components, we collect their latency by probing the model using DeepSpeed [30] while varying (1) the number of diffusion steps (5020), and (2) the span of cross-frame attention, i.e. self (1-frame), sparse (2-frames) or dense (8-frames). Based on the latency measurements aggregated over relevant groups of operations as reported in Fig. 2 we make the following observations:
@@
As demonstrated in Fig. 2 (a), inversion-based pipelines are susceptible to memory operations. Not only that the latency due to memory access dominates the overall latency in all the tested settings, but it also scales exponentially with the span (i.e., #frames) of cross-frame attention.
@@
When ruling out memory and considering computations only, as in Fig. 2 (b) and Fig. 2 (c), we appreciate that attention-related operations are as expensive as (or often more) than all other network operations combined (i.e. convolutions, normalizations, MLPs etc.) This latter finding is consistent both during inversion and generation phase.
@@
Using more-powerful cross-frame attentions (i.e., dense) that increases the temporal stability of generations [49], comes at a high latency cost as observed in Fig. 2 (b) and Fig. 2 (c).
@@ 3.1 Off-the-shelf acceleration
We explore whether the solutions available for image generation mitigate the key computational bottlenecks observed for video editing. More specifically, we study whether token merging [4], as an effective strategy to improve the attention cost for image generation [3], reduces the high memory and computational cost of video editing. Moreover, given the linear relation between the number of diffusion steps and memory costs of inversion-based editing to store the attention maps and activations [44, 33, 8, 46], we explore how using faster noise schedulers improves the video editing efficiency. In the followings, we describe how the off-the-shelf accelerations are adopted for video editing:
@@
Aiming to speed-up attention operations, we utilize ToMe [4, 3] to merge redundant tokens. More specifically, ToMe first splits the input self-attention tokens into source (src) and destination (dst) sets. Then, the similarity between the source token and the destination token is computed based on normalized cosine similarity as follows:
Finally, the top similarities are selected, and source tokens are piggybacked into the corresponding destination tokens by means of average pooling. This process reduces successive operations by removing merged tokens, as only destination and unmerged-source tokens (unm) are forwarded, at the cost of a slight information loss.
In our first round of experimentation, we observed the potential of ToMe in speeding up video editing, as a simple reduction of in key-value tokens reduces the model latency by a factor of 4. Nevertheless, its naive application completely breaks generation quality, as testified in Fig. 3 (d). However, we could restore an outstanding generation quality by applying two simple yet crucial tricks. Specifically, we i) pair the token locations (both dst and unm), for the same network layer and diffusion timestep, between inversion and generation (see ToMe-paired Fig. 3(e)), and ii) we resample the destination token locations for every frame (Fig. 3 (f)). We refer the reader to the supplementary material for more details about these ToMe optimizations.
@@
Although we observe that the commonly-used DDIM scheduler [43] suffers considerable quality degradation when reducing the number of diffusion steps (without any finetuning), we notice this is not the case with a more modern DPM++ [27], that can operate with as few as 20 iterations. As illustrated in the example reported in Fig. 3 (c), this replacement does not affect the quality of the editing, yet it impacts the generation latency tremendously.
In the remainder of the paper, we refer to models optimized with these off-the-shelf accelerations as optimized, and will serve as strong baselines to our proposed Object Centric Diffusion.
@@ 4 Object-Centric Diffusion
Although off-the-shelf optimizations can notably improve latency, further reductions can be achieved by exploiting the object-centric nature of video editing tasks. As such, following our key assumption regarding the significance of foreground objects for many video editing applications, we propose two add-ons for diffusion models to channel their computational capacity towards foreground regions: i) a diffusion sampling process that performs more sampling steps around the foreground regions as described in Sec. 4.1, and ii) a token reduction technique to minimize the background tokens, complemented by an effective spatiotemporal merging strategy outlined in Sec. 4.2. For both these contributions, we assume we have a foreground mask , highlighting in every frame the locations of objects to be edited. Such mask can be obtained by means of a segmentation model [21], visual saliency [34, 47] or by cross-attention with text-prompts.
@@ 4.1 Object-Centric Sampling
The process of sampling in diffusion models entails gradually transitioning from the initial noise to the sample derived from the real data distribution through denoising steps . Each denoising step involves running a computationally-heavy denoising UNet [39], where high-quality generations demand a large number of sampling steps. The state-of-the-art methods rely on efficient noise schedulers [26, 27] and step distillation [40, 25, 23] to minimize the number of required sampling steps.
Complementing the above, we introduce an efficient object-centric diffusion sampling technique tailored for editing tasks, prioritizing high-quality generation specifically in designated foreground regions.
As described in Algorithm 1, we divide the latent variable according to the foreground mask (e.g. a segmentation mask commonly available in video datasets or from a detector) into foreground and background latents denoted by and . Instead of running the same number of sampling steps for the foreground and background regions, we reduce the sampling steps on the background based on the hyperparameter . To avoid inconsistent generation across foreground and background, we merge the denoised latents at a certain sampling step specified by the hyperparameter . We empirically observed that allocating of the sampling steps to this blending stage is usually sufficient to generate spatially and temporally consistent frames.
We note that for localized editing tasks, e.g. shape and attribute manipulations of objects, the background sampling stage can be completely skipped, which results in even faster generation and higher fidelity reconstruction of background regions, as demonstrated in Fig. 7.
Given potentially irregular shape of the foreground masks, sampling of foreground and background regions involves running the UNet on sparse inputs. This requires using efficient sparse convolution and transformers as implemented in [22, 37, 45, 10]. To avoid using custom sparse operations in our implementation, we pad the mask to have a rectangular shape then extract a full crop around the foreground object, and use dense computations to denoise this crop. This simple implementation comes at the cost of not fully-utilizing the sparsity of foreground objects, that can be further improved by integrating the sparse engine implementation from [22]. As demonstrated in our experiments, a naive implementation still results in significant speed-ups, especially for smaller foreground regions.
4.2 Object-Centric 3D Token Merging

### v2

@@ Abstract
Diffusion-based video editing have reached impressive quality and can transform either the global style, local structure, and attributes of given video inputs, following textual edit prompts. However, such solutions typically incur heavy memory and computational costs to generate temporally-coherent frames, either in the form of diffusion inversion and/or cross-frame attention. In this paper, we conduct an analysis of such inefficiencies, and suggest simple yet effective modifications that allow significant speed-ups whilst maintaining quality. Moreover, we introduce Object-Centric Diffusion, to fix generation artifacts and further reduce latency by allocating more computations towards foreground edited regions, arguably more important for perceptual quality. We achieve this by two novel proposals: i) Object-Centric Sampling, decoupling the diffusion steps spent on salient or background regions and spending most on the former, and ii) Object-Centric Token Merging, which reduces cost of cross-frame attention by fusing redundant tokens in unimportant background regions. Both techniques are readily applicable to a given video editing model without retraining, and can drastically reduce its memory and computational cost. We evaluate our proposals on inversion-based and control-signal-based editing pipelines, and show a latency reduction up to 10 for a comparable synthesis quality. Project page: qualcomm-ai-research.github.io/object-centric-diffusion.
Keywords:
Video Editing Efficiency Object-Centric Diffusion1 Introduction
@@
Diffusion models [42, 11] stand as the fundamental pillar of contemporary generative AI approaches [37, 12, 17]. Their success primarily stems from their unparalleled diversity and synthesis quality, surpassing the capabilities of earlier versions of generative models [18, 8]. On top of that, recent methods such as Latent Diffusion Models (LDMs) [37] exhibit scalability to high-resolution inputs and adaptability to different domains without requiring retraining or finetuning, especially when trained on large-scale datasets [40]. These traits catalyzed impactful adaptations of pretrained LDMs across a spectrum of applications, including text-guided image generation [37], editing [1], inpainting [37], as well as video generation [2] and editing [44, 32, 47].
@@
Nevertheless, diffusion models come with various trade-offs. One immediate drawback is their inefficient sampling process, which involves iterating a denoising neural network across numerous diffusion steps. Despite the availability of techniques like step distillation [26, 39] and accelerated samplers [24, 25, 42] that expedite image synthesis, efficient sampling solutions for video generation are still lacking. Besides, special attention must be paid to maintaining temporal coherency among frames when creating or modifying a video, in order to avoid flickering artifacts or lack of correlation between frames. To this aim, approaches like diffusion inversion [42, 32] and cross-frame self-attention [32, 47] have been introduced, but all at the cost of further increasing the computational load.
@@
This paper centers on video editing models and presents novel solutions to enhance their efficiency. We first examine the current video editing frameworks, identifying the key elements that increase their latency. These encompass memory overheads, such as those associated with attention-based guidance from diffusion inversion, as well as computational bottlenecks, including excessive cross-frame attention and an unnecessarily high number of sampling steps. We show that significant improvements can be achieved by adopting off-the-shelf optimizations namely efficient samplers and leveraging token reduction techniques in attention layers such as token merging (ToMe) [3, 4].
@@
Additionally, we argue that video editing users may be particularly sensitive to the quality of edited foreground objects, as opposed to slight degradations in background regions. Consequently, we propose two new and efficient techniques that harness this object-centric aspect of editing applications. Our first solution, Object-Centric Sampling, involves separating the diffusion process between edited objects and background regions. This strategy enables the model to focus most of its diffusion steps on the foreground areas, whilst generating unedited regions with considerably fewer steps, yet maintaining faithful reconstructions in all areas. With our second solution, Object-Centric ToMe, we incentivise each token reduction layer to merge more tokens within less crucial background areas and exploit temporal redundancies that are abundant in videos. The combination of such techniques, coined Object-Centric Diffusion (OCD), can seamlessly be applied to any video editing method without retraining or finetuning.
@@
As we demonstrate in extensive experiments, the combined implementation of these object-centric solutions enhances the quality of video editing in both salient foreground object and background, while considerably reducing the generation cost. We accelerate both inversion-based and ControlNet-based video editing models, speeding them up by a factor of 10 and 6 respectively while also decreasing memory consumption up to 17 for a comparable quality (see Fig. 1).
@@
We introduce Object-Centric ToMe, which reduces number of cross-frame attention tokens by encouraging their fusion in background regions.
@@ 2 Related work
There has been a surge in text-to-video generation methods using diffusion models. The early works inflate the image generation architectures by replacing most operations, i.e. convolutions and transformers, with their spatio-temporal counterparts [13, 10, 14, 41, 48]. Despite their high temporal consistency, these substantial model modifications require extensive model training on large collection of captioned videos.
To avoid extensive model trainings, recent works adopt off-the-shelf image generation models for video editing in one-shot and zero-shot settings. One-shot methods rely on test-time model adaption on the editing sample, which is unfeasible for real-time applications [44, 6]. Zero-shot methods integrate training-free techniques into the image generation models to ensure temporal consistency across frames [32, 47, 15, 7, 23] most commonly by: i) providing strong structural conditioning from the input video [6, 47, 5] inspired by ControlNet [46]; ii) injecting the activations and attention maps extracted by diffusion inversion [42, 43] of the input video into the generation [32, 7]; iii) replacing the self-attention operations in the architecture with temporal counterparts operating on neighboring frames [44, 32, 47]. Introducing cross-frame attention greatly increases the temporal consistency, especially when involving more and more frames in the operation [47]. Despite their effectiveness, all these solutions come with additional computational costs, which are currently underexplored in the literature.
@@
Due to their sequential sampling, diffusion models are computationally very expensive. Several studies analyze the denoising U-Net architecture to enhance its runtime efficiency, achieved either through model distillation [16] or by eliminating redundant modules [22]. Other ways to obtain efficiency gains include enhanced noise schedulers [24, 25] and step distillation techniques [39, 22, 26]. Moreover, the work in [20] enables efficiency in image editing applications by caching representations of regions left unedited, yet it cannot handle videos and it strictly requires human-in-the-loop to operate. The most relevant to our work is the Token Merging technique presented in [3] that demonstrated advantages for image generation [4]. Although very proficient, we observe it does not readily work when deployed in video editing settings. For this reason, we optimize this technique for the task at hand, such that it exploits redundancy of tokens across frames and directing its token fusing towards background regions. Finally, Token Merging was applied to video editing use cases in VidToMe [21], where its introduction is aimed at improving temporal consistency of generated frames. Differently, we tailor it to improve efficiency without sacrificing fidelity in edited objects, resulting in much faster generations, as we shall demonstrate in our experimentation.
@@ 3 Efficiency Bottlenecks
To investigate the main latency bottlenecks in video editing pipelines, we run a benchmark analysis on a representative pipeline, derived from FateZero [32]. More specifically, we are interested in the role of components such as diffusion inversion and cross-frame attention, as these techniques are crucial for the task and ubiquitous in literature [44, 47, 32, 7, 21]. To study their impact, we perform text-guided edits on 8-frame clips while varying (1) the number of diffusion steps (5020), and (2) the span of cross-frame attention, i.e. self (1-frame), sparse (2-frames) or dense (8-frames). In doing so, we probe the model’s latency using DeepSpeed [28]. Based on the latency measurements aggregated over relevant groups of operations as reported in Fig. 2, we make the following observations:
@@
As testified by Fig. 2 (a), inversion-based pipelines are susceptible to memory operations. Not only the latency due to memory access dominates the overall latency in all the tested settings, but it also scales exponentially with the span (i.e., #frames) of cross-frame attention.
@@
When ruling out memory and considering computations only, as in Fig. 2 (b) and Fig. 2 (c), we appreciate that attention-related operations are as expensive as (or often more) than all other network operations combined (i.e. convolutions, normalizations, MLPs etc.). This latter finding is consistent both in inversion and generation phases.
@@
More-powerful cross-frame attention operations (i.e., dense), that increase the temporal stability of generations [47], come with a high latency cost, as observed in Fig. 2 (b) and Fig. 2 (c).
@@ 3.1 Off-the-shelf acceleration
We explore whether the solutions available for image generation mitigate the key computational bottlenecks observed for video editing. More specifically, we study whether token merging [3], as an effective strategy to improve the attention cost for image generation [4], reduces the high memory and computational cost of video editing. Moreover, given the linear relation between the number of diffusion steps and memory costs of inversion-based editing to store the attention maps and activations [43, 32, 7, 44], we explore how using faster noise schedulers improves the video editing efficiency. In what follows, we describe how the off-the-shelf accelerations are adopted for video editing:
@@
Aiming to speed-up attention operations, we utilize Token Merging (ToMe) [3, 4] to merge redundant tokens. More specifically, ToMe first splits the input self-attention tokens into source (src) and destination (dst) sets. Then, the similarity between the source token and the destination token is computed based on normalized cosine similarity as follows:
Finally, the top similarities are selected, and source tokens are piggybacked into the corresponding destination tokens by means of average pooling. This process lightens successive operations as only destination and unmerged-source tokens (unm) are forwarded, at the cost of a slight information loss.
We observed the potential of ToMe in speeding up video editing, as a simple reduction of in key-value tokens reduces the model latency by a factor of 4. Yet, its naive application completely breaks generation quality, as testified in Fig. 3 (d). However, we could restore an outstanding generation quality by applying two simple yet crucial tricks. Specifically, we i) pair the token locations (both dst and unm), for the same network layer and diffusion timestep, between inversion and generation (see ToMe-paired Fig. 3(e)), and ii) we resample the destination token locations for every frame (Fig. 3 (f)). We refer the reader to the supplementary material for more details about these ToMe optimizations.
@@
Although we observe that the commonly-used DDIM scheduler [42] suffers considerable quality degradation when reducing the number of diffusion steps (without any finetuning),
we notice this is not the case with a more modern DPM++ [25], that can operate with as few as 20 iterations.
As illustrated in the example reported in Fig. 3 (c), this replacement does not affect the quality of the editing, yet it impacts the generation latency tremendously.
We will refer to models optimized with these off-the-shelf accelerations as optimized, and will serve as strong baselines to our proposed Object Centric Diffusion.
Such a baseline is extremely fast, yet prone to sporadic generation artifacts (Fig. 3 (f), Fig. 7), that we shall fix with the following object-centric solutions.
@@ 4 Object-Centric Diffusion
Although off-the-shelf optimizations can notably improve latency, they occasionally introduce some editing artifacts in foreground objects, that we solve by exploiting the object-centric nature of video editing tasks, also enabling further reductions in latency. Following our assumption regarding the significance of foreground objects for video editing applications, we propose two add-ons for diffusion models to channel their computational capacity towards foreground regions, and we describe them in Sec. 4.1 and in Sec. 4.2. We will assume access to a foreground mask , highlighting in every frame the locations of objects to be edited. Such mask can be obtained, for instance, with a pretrained segmentation model [19], visual saliency [33, 45] or by cross-attention with text-prompts.
@@ 4.1 Object-Centric Sampling
The process of sampling in diffusion models entails gradually transitioning from the initial noise to a sample from the real data distribution, obtained through a series of denoising steps . Each step involves running a computationally-heavy denoiser (e.g. UNet [38]), and high-quality generations demand a large number of sampling steps. We hereby introduce an efficient object-centric diffusion sampling technique tailored for editing tasks, prioritizing high-quality generation specifically in designated foreground regions.
Our Object-Centric Sampling scheme is described in Algorithm 1. Specifically, we first consider the foreground mask and split the latent variable into foreground and background latents denoted by and . This operation is performed by a generic gather function, splitting representation tokens or feature map pixels into two disjoint sets. Then, instead of performing a single diffusion process, we disentangle the generation of foreground and background regions: such a decoupled scheme allows us to reduce the sampling steps on the latter based on an hyperparameter . Once foreground and background latents have been generated, we rely on a scatter operation to recompose a full feature map. Nevertheless, we observed that performing this recomposition step at the end of the whole diffusion process (i.e. merging variables and ) typically results into weak editing performance, in the form of color distribution-shift and boundary artifacts between foreground and background regions. We therefore introduce two further solutions to this issue. First, we introduce a normalization step (inspired by batch normalization) applying an affine transformation to the foreground latents, shifting and scaling them to match the mean and standard deviation of background representations. Even more importantly, we move the scattering of foreground and background latents within the sampling process, at a given timestep which is controlled by a hyperparameter . This choice introduces some diffusion timesteps , that take care of smoothing out such artifacts and seamlessly blend foreground and background latents. We empirically observed that allocating (i.e. ) of the sampling steps to this blending stage is usually sufficient to generate spatially and temporally consistent frames. We also note that for localized editing tasks, e.g. shape and attribute manipulations of objects, the background sampling stage can be completely skipped, which results in even faster generation and higher fidelity reconstruction of background regions, as demonstrated in Fig. 7.
4.2 Object-Centric Token Merging